The diameter of each wheel of a bicycle is 28 inches. If you are traveling a speed of 10 miles per hour on this bicycle, through how many revolutions per minute are the wheels turning?
Answer by nerdybill(7384)

The diameter of each wheel of a bicycle is 28 inches. If you are traveling a speed of 10 miles per hour on this bicycle, through how many revolutions per minute are the wheels turning?
.
convert 10 miles into inches:
10 miles = 10*5280 = 52800 feet = 633600 inches
.
circumference of wheel:
(pi)d = (3.14)28 = 87.92 inches per revolution
.
633600/87.92 =7206.55 revolutions per hour
convert to revolutions per min:
7206.55 rev/hr * 1/60 hr/min
120.1 revolutions per minute